You are considering purchasing a CNC machine which costs ​$200,000. This machine will have an estimated service life of 9 years with a net​ after-tax salvage value of ​$20,000. Its annual after-tax operating and maintenance costs are estimated to be ​$58,000. To expect an 19​% rate of return on​ investment, what would be the required minimum annual​ after-tax revenues?

The required minimum annual​ after-tax revenues would be ​$____thousand. ​(Round to one decimal​ place.)
Answer #1

Let the minimum annual after-tax revenue per year be R. Then in break-even,

PW of annual revenue = Initial cost + PW of annual costs

R x P/A(19%, 9) = 200,000 + 40,000 x P/A(19%, 9) + 20,000 x P/F(19%, 9)

P/A(19%, 9) x (R - 58,000) = 200,000 + 20,000 x 0.2090

4.1633 x (R - 58,000) = 200,000 + 4,180

4.1633 x (R - 58,000) = 204,180

R - 58,000 = 49,042.83

R = 107,042.83